What is the volume of 1kg of nitrogen gas (diatomic) at 24∘C and .98atm?



1
Expert's answer
2021-05-05T07:12:10-0400

"PV = nRT"


"V= \frac{nRT}{P}"



"n= moles= \frac{1000}{28}=35.71moles"


"V= \frac{35.71×0.08206×297}{98}=8.88"